Fermi National Laboratory
 

Accelerator Update for April 2, 2008 - April 4, 2008

 

Wednesday April 2

The day shift began with Operations monitoring the Tevatron (TeV) store 6021 (140.66E30), with the Antiproton Source conducting a study, with the Recycler, the Main Injector, Linac, and Booster all off and in access, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, with NuMI replacing its bad quadrupole magnet (the main reason for the shutdown), and with the SciBooNE / MiniBooNE (BNB) experiments off, not because they have a problem or are doing work, but because all the other machines are off.

The Pbar study ended at 8:33 AM.

A kicker expert began work on the MI-10 kicker at 8:38 AM.

Water group technicians reported at 9:07 AM that they had completed the MI-60 heat exchanger (HE) flush.

Since the MiniBooNE experimenters were feeling a little left out, at 9:43 AM they decided they would access their hall. Operators prepared the area for access.

Linac experts reported at 11:12 AM that they had replaced the flow turbines for three Klystron RF stations (KRF1, 4, and 7).

At 11:42 AM, Booster experts reported that a kicker filament transformer had failed on MKS08 and that they were beginning repairs.

Linac came out of access at 12:37 PM.

At 12:58 PM, a Linac expert reported that the surge resisters for LRF3 had failed. They began replacing the station’s thyratron.

Booster experts reported at 1:17 PM that they had repaired MKS08.

Linac experts reported at 2:14 PM that they had completed their repair on LRF3, but now LRF5 was experiencing some problems. They began a LRF5 study.

Operations established beam to the Linac dump at 2:28 PM.

Linac experts completed their LRF5 study at 4:09 PM.

Booster came out of access and Operations established beam to the Booster dump at 4:54 PM.

A Booster vacuum pump (IPRF23) tripped off, which caused BRF13 and 14 to trip off. Everything reset.

Pbar experts completed installing their rebuilt lithium lens, operators search and secured the rings, and an expert began pulsing the lens at 5:43 PM.

At 6:55 PM, KRF3 tripped off for the first of two times in ten minutes. It reset.

A NuMI expert reported at 8:09 PM that the new QD111 magnet had been installed, the alignment crew was on hand, and the vacuum technicians would begin hooking up the magnet soon.

At 8:38 PM, LRF2 tripped off. It reset.

At 9:23 PM, a Booster vacuum pump (B:IP61) tripped off. It reset.

At 10:18 PM, Operations reported that a TeV sector D1 wet engine alarmed on a high cryostat vacuum indication.

Thursday April 3

The midnight shift began with Operations monitoring store 6021 (46.65E30), and with NuMI off due to a magnet replacement.

At 12:40 AM, LRF3 tripped off. It reset.

At 12:45 AM, KRF4 tripped off. It reset.

At 4:12 AM, BRF8 tripped off. It reset.

At 5:23 AM, KRF6 tripped off. It reset.

At 5:37 AM, KRF1 tripped off. It reset.

At 7 AM, a TeV Collider A0 Proton Abort Kicker (C:PAK4) tripped off on a Personnel Interlock fault. Operations contacted an expert.

Operations prepared Booster for an access at 7:01 AM.

At 8:12 AM, KRF7 tripped off. It reset.

At 8:47 AM, a Pbar Debuncher vacuum pump (D:IP403) tripped off. It reset.

A TeV expert reported at 9:02 AM that to compensate for PAK4 being off he had created a four-bump position change to move the A0 beam so it could be aborted with fewer kickers.

At 9:33 AM, KRF1 tripped off. It reset.

Operations reported at 9:41 AM that the PAK4 kicker had been repaired. The four-bump was removed.

At 9:49 AM, KRF4 tripped off. Operators had some trouble resetting this station, but it did reset.

Operations prepared Linac for an access at 10:08 AM.

Operations search and secured the Pelletron at 10:26 AM.

Linac came out of access and Operations tried to turn the accelerator on at 10:50 AM, but the Klystron vernier would not come on.

Operations reported at 10:59 AM that the NuMI magnet and its power supply QD111 were electrically sound and ready to operate.

Booster came out of access at 11:18 AM.

Operations prepared Linac for access at 11:20 AM.

A NuMI expert reported at 11:55 AM that QD111 was aligned.

Linac came out of access and Operations started sending beam to the Booster dump at noon. Booster experts tuned the beam.

At 12:53 PM, a Booster expert reported that a kicker (MKS08) was failing. They were going to prepare a new coil and then make repairs.

At 12:54 PM, KRF4 tripped off. It reset.

At 12:59 PM, RF experts tried to start the pond water pump for MIRF1, but it would not run. They contacted FESS personnel to help with the pump.

At 1:56 PM, LRF4 tripped off for the first of two times in about 40 minutes. It reset.

At 2:24 PM, BRF7 tripped off. It reset.

At 3:02 PM, BRF3 tripped off. It reset.

At 3:18 PM, KRF4 tripped off. Experts investigated and reset the station.

MI came out of access at 4 PM.

Operations established beam to BNB at 4:22 PM.

Operations established beam to NuMI at 5:09 PM.

Operations established beam to Pbar at 5:48 PM.

Pbar experts began an AP1 transfer line trajectory study and target tuning study at 6:06 PM.

At 6:16 PM, Operations reported that the TeV orbit monitor feedback was not updating. An operator restarted the DCE and that restored the orbit feedback.

Operations raised NuMI’s beam intensity to 6 turns at 6:41 PM.

A Recycler expert began a study at 7:15 PM.

CDF and D0 turned off their detector’s high voltages in preparation for a TeV end-of-store study at 7:38 PM.

At 7:56 PM, a Booster septa power supply (MP02) tripped off for the first of three times in four minutes on a ground fault indication. It reset.

Operations terminated store 6021 at 8:21 PM. Cryo system experts began work on the TeV A2 wet engine.

Operations prepared D0 for an access at 8:34 PM.

At 9:24 PM, three Pbar Debuncher beam valves closed (D:BV728, D:BV309, and D:BV210D). They all opened.

Cryo system experts completed their work on the A2 wet engine flywheel at 10:02 PM.

At 11:12 PM, the MI-10 area suffered a vacuum burst. It recovered.

At 11:58 PM, KRF2 tripped off. It reset.

Operations reported at midnight that the Recycler had stashed antiprotons twice during the evening shift.

Friday April 4

The midnight shift began with Operations monitoring the stash (50.9E10), with Pbar stacking (15.2E10), with D0 in access to fix a front-end problem, and with NuMI and BNB taking beam.

At 12:11 AM, a vacuum pump tripped off in the MI-10 area.

At 12:29 AM, MP02 tripped off on a ground fault. It resets, but continues to trip. Operations contacted experts.

At 3 AM, KRF2 tripped off. It reset.

At 4:06 AM, KRF1 tripped off on a LCW flow fault. It reset.

Operations resumed sending beam to NuMI, BNB, and Pbar at 5:04 AM.

At 5:16 AM, an MI RF center anode power supply was tripping off. MIRF12 was also tripping off and would not stay reset. Operators bypassed the RF station.

At 5:26 AM, BRF19 tripped off. It reset.

At 5:27 AM, Operations reported that the MP02 septa power supply was still tripping off. Experts were investigating.

At 6:13 AM, beam valve D:BV728 closed. It opened.

At 6:34 AM, beam valves D:BV728, D:BV309, and D:BV210D all closed and M:USWSTA tripped off. Everything opened and reset.

Booster experts reported at 7:16 AM that MP02 was faulting on every pulse earlier this morning. Then it began faulting about every 30 minutes. While watching two newly installed voltage monitors, the experts noticed that the ground voltage bounced up to 40 volts during the ground faults. They added a more sensitive current monitor to follow the ground voltage bounce. They suspected a possible load problem due to one ground fault that was followed by a MP02 vacuum trip. They reported that they would install more diagnostic equipment to help them find where the ground fault current was going.

At 7:47 AM, BRF13 tripped off. It reset.

An RF expert reported at 7:50 AM that he had brought MIRF12 back online.

Operations reported at 7:58 AM that the Recycler had stashed once during the midnight shift.

The Plans for Friday and the Weekend

The plans for today is to allow D0 to fix its front-end problems and for Booster experts to resolve its septa problems. The Run Coordinator would like to put the TeV into shot setup by noon. For the weekend he wants to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/04/2008    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory